# Drinfeld rational fractions for affine Kac–Moody quantum symmetric pairs

### v2

@@ 4.4. Proof of the factorization theorem
(4.52) $(1-c_{0}c_{1}q^{2}z^{2})\boldsymbol{\Theta}(z)\equiv$ $\ \frac{(K-CK^{-1}z^{2})\boldsymbol{\phi}(-q^{-2}z^{-1})}{q-q^{-1}}$
(4.53) $\ +c_{0}z\boldsymbol{\phi}(-q^{-2}z^{-1})[\mathbb{A},(1-q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)x_{1}^{-}]\quad\mod{U_{+}}[\negthinspace[z]\negthinspace].$
@@
(4.57) $\mbox{LHS}=-q^{-2}(z^{-1}x^{-}_{1}+x^{-}_{0})\boldsymbol{x}^{-}_{\leq 0}(-q^{-2}z^{-1})+\boldsymbol{x}^{-}_{\leq 0}(-q^{-2}z^{-1})(q^{-4}z^{-1}x^{-}_{1}+x^{-}_{0}).$
@@
$(-q^{2})^{s}\left(-q^{-2}(-q^{2}x^{-}_{1}x^{-}_{-s-1}+x^{-}_{0}x^{-}_{-s})+(-q^{-2}x^{-}_{-s-1}x^{-}_{1}+x^{-}_{-s}x^{-}_{0})\right)$
if $s\geq 0$ and $-q^{-2}[x^{-}_{1},x^{-}_{0}]_{q^{-2}}$ if $s=-1$. In both cases, the expression vanishes thanks to relation (2.8), with $k=0$ and $l=-s-1$.
@@
(4.66) $\equiv$ $\ \frac{(K-CK^{-1}z^{2})\boldsymbol{\phi}(-q^{-2}z^{-1})+(Cz^{2}-1)}{c_{1}^{-1}Cz^{2}(q-q^{-1})}.$
Moreover, by Lemma (4.6), we have
(4.67) $(z^{-1}A_{-1}^{(-1)}-q^{-2}A_{0}^{(-1)})\boldsymbol{A}^{(1)}_{+}(z)=$ $\ (z^{-1}A_{-1}^{(-1)}-q^{-2}A_{0}^{(-1)})\boldsymbol{\phi}(-q^{-2}z^{-1})\mathbb{A}$
(4.68) $=$ $\ -q^{-2}(x_{1}^{-}z^{-1}+x_{0}^{-})\boldsymbol{\phi}(-q^{-2}z^{-1})\mathbb{A}$
(4.69) $=$ $\ -\boldsymbol{\phi}(-q^{-2}z^{-1})(q^{-4}x_{1}^{-}z^{-1}+x_{0}^{-})\mathbb{A},$
@@
$\boldsymbol{A}^{(1)}_{+}(z)(-q^{-2}z^{-1}A_{-1}^{(-1)}+A_{0}^{(-1)})=\boldsymbol{\phi}(-q^{-2}z^{-1})\mathbb{A}(q^{-4}x^{-}_{1}z^{-1}+x^{-}_{0}).$
@@
(4.71) $=$ $\ q^{-4}z^{-1}\boldsymbol{\phi}(-q^{-2}z^{-1})[\mathbb{A},(1-q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)x_{1}^{-}].$
@@
(4.72) $(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)\mathbb{A}=-c_{1}q^{2}(1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\boldsymbol{x}^{+}_{\geq 0}(-c_{0}c_{1}q^{2}z)$
@@
(4.74) $-(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(-q^{-4}z^{-1})=$ $\ (q^{6}-q^{10})x^{+}_{-2}z^{2}$
(4.75) $(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)(-q^{2}x^{+}_{0}+q^{6}x^{+}_{-1}z)=$ $\ q^{10}x^{+}_{-2}z^{2}-q^{2}x^{+}_{0}.$
@@
(4.76) $(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)M=q^{6}x^{+}_{-2}z^{2}-q^{2}x^{+}_{0}=(1-q^{4}\operatorname{ad}_{\overline{h}_{-1}}^{2}z^{2})(-q^{2}x^{+}_{0})$
@@
(4.77) $(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)\Omega^{-1}M=-q^{2}\frac{1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z}{1+c_{0}c_{1}q^{2}\operatorname{ad}_{\overline{h}_{1}}z}x^{+}_{0}.$
@@
(4.79) $(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)\mathbb{A}=$ $\ c_{1}(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)\Omega^{-1}M$
(4.80) $=$ $\ -q^{2}\frac{1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z}{1+c_{0}c_{1}q^{2}\operatorname{ad}_{\overline{h}_{1}}z}x^{+}_{0}$
(4.81) $=$ $\ -c_{1}q^{2}(1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\boldsymbol{x}^{+}_{\geq 0}(-c_{0}c_{1}q^{2}z),$
@@
(4.82) $\boldsymbol{\Theta}(z)\equiv\frac{1-Cz^{2}}{(q-q^{-1})(1-q^{-2}Cz^{2})}\boldsymbol{\phi}(-q^{-2}z^{-1})\boldsymbol{\psi}(-Cq^{-2}z)\mod{U_{+}}[\negthinspace[z]\negthinspace].$
@@
(4.83) $c_{0}z\boldsymbol{\phi}(-q^{-2}z^{-1})[\mathbb{A},(1\ -$ $\ q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)x_{1}^{-}]=c_{0}z\boldsymbol{\phi}(-q^{-2}z^{-1})[(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)\mathbb{A},x_{1}^{-}]$
(4.84) $=$ $\ -Cq^{-2}z\boldsymbol{\phi}(-q^{-2}z^{-1})[(1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\boldsymbol{x}^{+}_{\geq 0}(-Cq^{-2}z),x^{-}_{1}]$
(4.85) $=$ $\ -Cq^{-2}z\boldsymbol{\phi}(-q^{-2}z^{-1})[\boldsymbol{x}^{+}_{\geq 0}(-Cq^{-2}z),(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x^{-}_{1}]$
(4.86) $=$ $\ \frac{-Cq^{-2}z\boldsymbol{\phi}(-q^{-2}z^{-1})}{q-q^{-1}}\big{(}(-Cq^{-2}z)^{-1}\boldsymbol{\psi}_{\geq 1}(-Cq^{-2}z)$
(4.87) $\ +q^{2}z\boldsymbol{\psi}(-Cq^{-2}z)-q^{2}zK^{-1}\big{)}$
(4.88) $=$ $\ \frac{\boldsymbol{\phi}(-q^{-2}z^{-1})}{q-q^{-1}}\big{(}(1-Cz^{2})\boldsymbol{\psi}(-Cq^{-2}z)-K+Cz^{2}K^{-1}\big{)}.$
@@
(4.90) $\boldsymbol{\grave{\Theta}}(z)\equiv\boldsymbol{\phi}(-q^{-2}z^{-1})\boldsymbol{\psi}(-Cq^{-2}z){\mod U_{+}}[\negthinspace[z]\negthinspace].$
@@ 5.1. Spectra of restricted representations
The $\mathcal{O}_{q}$-module $\eta_{\mathbf{c},0}^{*}(V)$ has a decomposition into generalized $\mathcal{H}_{\mathcal{O}_{q}}$-eigenspaces. Let $w$ be generalized $\mathcal{H}_{\mathcal{O}_{q}}$-eigenvector and $\overline{w}$ its image in $\operatorname{gr}(V)$. Corollary 4.11 implies that $\overline{w}$ is an $\mathcal{H}$-eigenvector, and that the $\mathcal{H}_{\mathcal{O}_{q}}$-eigenvalues of $w$ can be computed using the $\mathcal{H}$-eigenvalues of $\overline{w}$. More precisely, in terms of generating series, the eigenvalue of $\boldsymbol{\grave{\Theta}}(z)$ on $w$ is the same as the $\boldsymbol{\phi}(-q^{-2}z^{-1})\boldsymbol{\psi}(-Cq^{-2}z)$-eigenvalue on $\overline{w}$. This fact has many interesting consequences.
@@ 5.2. Frenkel–Reshetikhin-type theorem
(5.3) $P(z)=\gamma_{P}z^{\deg P}P^{\dagger}(C^{-1}z^{-1}),\qquad P(-z)=\gamma_{P}z^{\deg P}P^{*}(-z^{-1}),$
@@
$\mathbf{Q}(z)={Q(-Cq^{-2}z)R^{*}(-q^{2}z)},\quad\mathbf{Q}^{\dagger}(z)={R(-Cq^{-2}z)Q^{*}(-q^{2}z)}.$
@@
$\boldsymbol{D}_{w}(z)=\boldsymbol{d}_{w}^{+}(-Cq^{-2}z)\boldsymbol{d}_{w}^{-}(-q^{-2}z^{-1}).$
@@
(5.5) $\boldsymbol{D}_{w}(z)=$ $\ \frac{Q^{*}(-q^{3}z)R^{*}(-qz)Q(-Cq^{-3}z)R(-Cq^{-1}z)}{Q(-Cq^{-1}z)R(-Cq^{-3}z)Q^{*}(-qz)R^{*}(-q^{3}z)}$
(5.6) $=$ $\ \left(\frac{Q(-Cq^{-3}z)R^{*}(-qz)}{Q(-Cq^{-1}z)R^{*}(-q^{3}z)}\right)\left(\frac{R(-Cq^{-1}z)Q^{*}(-q^{3}z)}{R(-Cq^{-3}z)Q^{*}(-qz)}\right),$
@@ 7.1. Coproduct formula for $\boldsymbol{A}_{+}(z)$
$\Delta(\overline{H}_{1})=1\otimes\overline{H}_{1}+\overline{H}_{1}\otimes 1+q^{2}(q-q^{-1})A_{0}\otimes x^{+}_{-1}-C(q-q^{-1})A_{-1}\otimes x^{+}_{0}.$
@@
(7.8) $[Q_{4},S^{r}_{1}]+[Q_{2},S^{r}_{2}]=$ $\ -CA_{-1}\otimes(q-q^{-1})[x^{+}_{0},(-q^{2})^{r}x^{-}_{-r}]+[\overline{H}_{1},A_{0}]\otimes(-q^{2})^{r}\phi_{-r}$
(7.9) $=$ $\ ([\overline{H}_{1},A_{0}]+CA_{-1})\otimes(-q^{2})^{r}\phi_{-r}=A_{1}\otimes(-q^{2})^{r}\phi_{-r}.$
@@
where $\widetilde{\xi}_{\mathbf{c},\mathbf{s}}$ is as in (2.23). From now on all the elements on the LHS of the tensor symbol are understood as images under $\xi_{\mathbf{c},\mathbf{s}}$. For example, $A_{0}\otimes(-q^{2})^{r}\phi_{-r}=\xi_{\mathbf{c},\mathbf{s}}(A_{0})\otimes(-q^{2})^{r}\phi_{-r}$, etc. In particular, any two elements on the LHS of the tensor symbol commute. Abbreviate
$\kappa=q^{2}(q-q^{-1})\big{(}\xi_{\mathbf{c},\mathbf{s}}(A_{0})-q^{-2}C\xi_{\mathbf{c},\mathbf{s}}(A_{-1})\operatorname{ad}_{\overline{h}_{1}}\big{)}$
@@
(7.12) $S^{r}_{4}=\sum_{\begin{subarray}{c}i+j+k=r;\\ i\geq 1;j,k\geq 0\end{subarray}}A_{k}\otimes\kappa\mathbbm{h}_{i-1}[x^{+}_{-1},(-q^{2})^{j}\phi_{-j}],$
where $\mathbbm{h}_{i}$ denotes the $i$-th degree complete homogeneous symmetric polynomial in $q^{2}\operatorname{ad}_{\overline{h}_{-1}}$ and $-q^{-2}C\operatorname{ad}_{\overline{h}_{1}}$, so that $\sum_{i=0}^{\infty}\mathbbm{h}_{i}z^{i}=\Omega^{-1},$ with $\Omega$ as in (4.32).
@@
(7.13) $[\overline{H}_{1},\mathbbm{h}_{i-1}]+C\mathbbm{h}_{i-2}\equiv$ $\ (q^{2}\operatorname{ad}_{\overline{h}_{-1}}-q^{-2}C\operatorname{ad}_{\overline{h}_{1}})\mathbbm{h}_{i-1}-(q^{2}\operatorname{ad}_{\overline{h}_{-1}})(-q^{-2}C\operatorname{ad}_{\overline{h}_{1}})\mathbbm{h}_{i-2}$
(7.14) $=$ $\ \mathbbm{h}_{i}.$
@@
(7.16) $\overline{\Delta}_{\mathbf{c},\mathbf{s}}(\boldsymbol{A}_{+}(z))\equiv$ $\ 1\otimes\boldsymbol{A}_{+}(z)+\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(-q^{-2}z^{-1})+\boldsymbol{A}_{+}(z)\otimes\kappa\Gamma,$
@@
$\Gamma=(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(1+q^{-2}C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(q^{2}-1)\boldsymbol{\phi}(-q^{-2}z^{-1})x^{+}_{-1}z.$
@@
$\sum_{r\geq 0}S^{r}_{1}=1\otimes\boldsymbol{A}_{+}(z),\qquad\sum_{r\geq 0}S^{r}_{2+3}=\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(-q^{-2}z^{-1}).$
@@
(7.17) $\sum_{r\geq 1}S^{r}_{4}z^{r}=$ $\ \boldsymbol{A}_{+}(z)\otimes z\kappa\Omega^{-1}[x^{+}_{-1},\boldsymbol{\phi}(-q^{-2}z^{-1})]=\boldsymbol{A}_{+}(z)\otimes\kappa\Gamma,$
@@
$\widetilde{\Gamma}=(1+q^{4}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(1+q^{-2}C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(q^{2}-1)x^{+}_{-1}z$
so that $\Gamma=\boldsymbol{\phi}(-q^{-2}z^{-1})\widetilde{\Gamma}$.

### v3

@@ 4.4. Proof of the factorization theorem
(4.49) $(1-q^{-2}Cz^{2})\boldsymbol{\Theta}(z)\equiv$ $\ \frac{(K-CK^{-1}z^{2})\boldsymbol{\phi}(z^{-1})}{q-q^{-1}}$
(4.50) $\ -q^{2}c_{0}z\boldsymbol{\phi}(z^{-1})[\mathbb{A},(1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x_{1}^{-}]\quad\mod{U_{+}}[\negthinspace[z]\negthinspace].$
@@
(4.54) $\mbox{LHS}=(z^{-1}x^{-}_{1}-q^{-2}x^{-}_{0})\boldsymbol{x}^{-}_{\leq 0}(z^{-1})+\boldsymbol{x}^{-}_{\leq 0}(z^{-1})(q^{-2}z^{-1}x^{-}_{1}+x^{-}_{0}).$
@@
$(x^{-}_{1}x^{-}_{-s-1}-q^{-2}x^{-}_{0}x^{-}_{-s})+(-q^{-2}x^{-}_{-s-1}x^{-}_{1}+x^{-}_{-s}x^{-}_{0})$
if $s\geq 0$ and $[x^{-}_{1},x^{-}_{0}]_{q^{-2}}$ if $s=-1$. In both cases, the expression vanishes thanks to relation (2.8), with $k=0$ and $l=-s-1$.
@@
(4.63) $\equiv$ $\ \frac{(K-CK^{-1}z^{2})\boldsymbol{\phi}(z^{-1})+(Cz^{2}-1)}{c_{1}^{-1}Cz^{2}(q-q^{-1})}.$
Moreover, by Lemma 4.6, we have
(4.64) $(z^{-1}A_{-1}^{(-1)}-q^{-2}A_{0}^{(-1)})\boldsymbol{A}^{(1)}_{+}(z)=$ $\ (z^{-1}A_{-1}^{(-1)}-q^{-2}A_{0}^{(-1)})\boldsymbol{\phi}(z^{-1})\mathbb{A}$
(4.65) $=$ $\ (x_{1}^{-}z^{-1}-q^{-2}x_{0}^{-})\boldsymbol{\phi}(z^{-1})\mathbb{A}$
(4.66) $=$ $\ \boldsymbol{\phi}(z^{-1})(q^{-2}x_{1}^{-}z^{-1}-x_{0}^{-})\mathbb{A},$
@@
$\boldsymbol{A}^{(1)}_{+}(z)(-q^{-2}z^{-1}A_{-1}^{(-1)}+A_{0}^{(-1)})=\boldsymbol{\phi}(z^{-1})\mathbb{A}(-q^{-2}x^{-}_{1}z^{-1}+x^{-}_{0}).$
@@
(4.68) $=$ $\ -q^{-2}z^{-1}\boldsymbol{\phi}(z^{-1})[\mathbb{A},(1+q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x_{1}^{-}].$
@@
(4.69) $(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\mathbb{A}=-c_{1}q^{2}(1-\operatorname{ad}_{\overline{h}_{-1}}z)\boldsymbol{x}^{+}_{\geq 0}(Cz)$
@@
(4.71) $-(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)(q^{2}-q^{-2})\boldsymbol{x}^{+}_{\leq-2}(q^{-2}z^{-1})=$ $\ (q^{2}-q^{6})x^{+}_{-2}z^{2}$
(4.72) $-(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)(q^{2}x^{+}_{0}+q^{4}x^{+}_{-1}z)=$ $\ q^{6}x^{+}_{-2}z^{2}-q^{2}x^{+}_{0}.$
@@
(4.73) $(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)M=q^{2}x^{+}_{-2}z^{2}-q^{2}x^{+}_{0}=(1-\operatorname{ad}_{\overline{h}_{-1}}^{2}z^{2})(-q^{2}x^{+}_{0})$
@@
(4.74) $(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\Omega^{-1}M=-q^{2}\frac{1-\operatorname{ad}_{\overline{h}_{-1}}z}{1-C\operatorname{ad}_{\overline{h}_{1}}z}x^{+}_{0}.$
@@
(4.76) $(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\mathbb{A}=$ $\ c_{1}(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\Omega^{-1}M$
(4.77) $=$ $\ -q^{2}c_{1}\frac{1-\operatorname{ad}_{\overline{h}_{-1}}z}{1-C\operatorname{ad}_{\overline{h}_{1}}z}x^{+}_{0}$
(4.78) $=$ $\ -c_{1}q^{2}(1-\operatorname{ad}_{\overline{h}_{-1}}z)\boldsymbol{x}^{+}_{\geq 0}(Cz),$
@@
(4.79) $\boldsymbol{\Theta}(z)\equiv\frac{1-Cz^{2}}{(q-q^{-1})(1-q^{-2}Cz^{2})}\boldsymbol{\phi}(z^{-1})\boldsymbol{\psi}(Cz)\mod{U_{+}}[\negthinspace[z]\negthinspace].$
@@
(4.80) $-q^{2}c_{0}z\boldsymbol{\phi}(z^{-1})[\mathbb{A},(1\ +$ $\ q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)x_{1}^{-}]=-q^{2}c_{0}z\boldsymbol{\phi}(z^{-1})[(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)\mathbb{A},x_{1}^{-}]$
(4.81) $=$ $\ Cz\boldsymbol{\phi}(z^{-1})[(1-\operatorname{ad}_{\overline{h}_{-1}}z)\boldsymbol{x}^{+}_{\geq 0}(Cz),x^{-}_{1}]$
(4.82) $=$ $\ Cz\boldsymbol{\phi}(z^{-1})[\boldsymbol{x}^{+}_{\geq 0}(Cz),(1+\operatorname{ad}_{\overline{h}_{-1}}z)x^{-}_{1}]$
(4.83) $=$ $\ \frac{Cz\boldsymbol{\phi}(z^{-1})}{q-q^{-1}}\big{(}(Cz)^{-1}\boldsymbol{\psi}_{\geq 1}(Cz)-z\boldsymbol{\psi}(Cz)+zK^{-1}\big{)}$
(4.84) $=$ $\ \frac{\boldsymbol{\phi}(z^{-1})}{q-q^{-1}}\big{(}(1-Cz^{2})\boldsymbol{\psi}(Cz)-K+Cz^{2}K^{-1}\big{)}.$
@@
(4.86) $\boldsymbol{\grave{\Theta}}(z)\equiv\boldsymbol{\phi}(z^{-1})\boldsymbol{\psi}(Cz){\mod U_{+}}[\negthinspace[z]\negthinspace].$
@@ 5.1. Spectra of restricted representations
The $\mathcal{O}_{q}$-module $\eta_{\mathbf{c},0}^{*}(V)$ has a decomposition into generalized $\mathcal{H}_{\mathcal{O}_{q}}$-eigenspaces. Let $w$ be generalized $\mathcal{H}_{\mathcal{O}_{q}}$-eigenvector and $\overline{w}$ its image in $\operatorname{gr}(V)$. Corollary 4.11 implies that $\overline{w}$ is an $\mathcal{H}$-eigenvector, and that the $\mathcal{H}_{\mathcal{O}_{q}}$-eigenvalues of $w$ can be computed using the $\mathcal{H}$-eigenvalues of $\overline{w}$. More precisely, in terms of generating series, the eigenvalue of $\boldsymbol{\grave{\Theta}}(z)$ on $w$ is the same as the $\boldsymbol{\phi}(z^{-1})\boldsymbol{\psi}(Cz)$-eigenvalue on $\overline{w}$. This fact has many interesting consequences.
@@ 5.2. Frenkel–Reshetikhin-type theorem
(5.3) $P(z)=\gamma_{P}z^{\deg P}P^{\dagger}(C^{-1}z^{-1}),\qquad P(z)=\gamma_{P}z^{\deg P}P^{*}(z^{-1}),$
@@
$\mathbf{Q}(z)={Q(Cz)R^{*}(z)},\quad\mathbf{Q}^{\dagger}(z)={R(Cz)Q^{*}(z)}.$
@@
$\boldsymbol{D}_{w}(z)=\boldsymbol{d}_{w}^{+}(Cz)\boldsymbol{d}_{w}^{-}(z^{-1}).$
@@
(5.5) $\boldsymbol{D}_{w}(z)=$ $\ \frac{Q^{*}(qz)R^{*}(q^{-1}z)Q(Cq^{-1}z)R(Cqz)}{Q(Cqz)R(Cq^{-1}z)Q^{*}(q^{-1}z)R^{*}(qz)}$
(5.6) $=$ $\ \left(\frac{Q(Cq^{-1}z)R^{*}(q^{-1}z)}{Q(Cqz)R^{*}(qz)}\right)\left(\frac{R(Cqz)Q^{*}(qz)}{R(Cq^{-1}z)Q^{*}(q^{-1}z)}\right),$
@@ 7.1. Coproduct formula for $\boldsymbol{A}_{+}(z)$
$\Delta(\overline{H}_{1})=1\otimes\overline{H}_{1}+\overline{H}_{1}\otimes 1-(q-q^{-1})A_{0}\otimes x^{+}_{-1}-C(q-q^{-1})A_{-1}\otimes x^{+}_{0}.$
@@
(7.8) $[Q_{4},S^{r}_{1}]+[Q_{2},S^{r}_{2}]=$ $\ -CA_{-1}\otimes(q-q^{-1})[x^{+}_{0},x^{-}_{-r}]+[\overline{H}_{1},A_{0}]\otimes\phi_{-r}$
(7.9) $=$ $\ ([\overline{H}_{1},A_{0}]+CA_{-1})\otimes\phi_{-r}=A_{1}\otimes\phi_{-r}.$
@@
where $\widetilde{\xi}_{\mathbf{c},\mathbf{s}}$ is as in (2.23). From now on all the elements on the LHS of the tensor symbol are understood as images under $\xi_{\mathbf{c},\mathbf{s}}$. For example, $A_{0}\otimes\phi_{-r}=\xi_{\mathbf{c},\mathbf{s}}(A_{0})\otimes\phi_{-r}$, etc. In particular, any two elements on the LHS of the tensor symbol commute. Abbreviate
$\kappa=-(q-q^{-1})\big{(}\xi_{\mathbf{c},\mathbf{s}}(A_{0})+C\xi_{\mathbf{c},\mathbf{s}}(A_{-1})\operatorname{ad}_{\overline{h}_{1}}\big{)}$
@@
(7.10) $S^{r}_{4}=\sum_{\begin{subarray}{c}i+j+k=r;\\ i\geq 1;j,k\geq 0\end{subarray}}A_{k}\otimes\kappa\mathbbm{h}_{i-1}[x^{+}_{-1},\phi_{-j}],$
where $\mathbbm{h}_{i}$ denotes the $i$-th degree complete homogeneous symmetric polynomial in $-\operatorname{ad}_{\overline{h}_{-1}}$ and $C\operatorname{ad}_{\overline{h}_{1}}$, so that $\sum_{i=0}^{\infty}\mathbbm{h}_{i}z^{i}=\Omega^{-1},$ with $\Omega$ as in (4.29).
@@
(7.11) $[\overline{H}_{1},\mathbbm{h}_{i-1}]+C\mathbbm{h}_{i-2}\equiv$ $\ (C\operatorname{ad}_{\overline{h}_{1}}-\operatorname{ad}_{\overline{h}_{-1}})\mathbbm{h}_{i-1}+C\mathbbm{h}_{i-2}-C\operatorname{ad}_{\overline{h}_{1}}\operatorname{ad}_{\overline{h}_{-1}}\mathbbm{h}_{i-2}=\mathbbm{h}_{i}.$
@@
(7.13) $\overline{\Delta}_{\mathbf{c},\mathbf{s}}(\boldsymbol{A}_{+}(z))\equiv$ $\ 1\otimes\boldsymbol{A}_{+}(z)+\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1})+\boldsymbol{A}_{+}(z)\otimes\kappa\Gamma,$
@@
$\Gamma=(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(1-C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(q^{2}-1)\boldsymbol{\phi}(z^{-1})x^{+}_{-1}z.$
@@
$\sum_{r\geq 0}S^{r}_{1}=1\otimes\boldsymbol{A}_{+}(z),\qquad\sum_{r\geq 0}S^{r}_{2+3}=\boldsymbol{A}_{+}(z)\otimes\boldsymbol{\phi}(z^{-1}).$
@@
(7.14) $\sum_{r\geq 1}S^{r}_{4}z^{r}=$ $\ \boldsymbol{A}_{+}(z)\otimes z\kappa\Omega^{-1}[x^{+}_{-1},\boldsymbol{\phi}(z^{-1})]=\boldsymbol{A}_{+}(z)\otimes\kappa\Gamma,$
@@
$\widetilde{\Gamma}=(1-q^{2}\operatorname{ad}_{\overline{h}_{-1}}z)^{-1}(1-C\operatorname{ad}_{\overline{h}_{1}}z)^{-1}(q^{2}-1)x^{+}_{-1}z$
so that $\Gamma=\boldsymbol{\phi}(z^{-1})\widetilde{\Gamma}$.